OpenStudy (anonymous):

please help. Aldo is a software salesman. His base salary is $2300 , and he makes an additional $120 for every copy of History is Fun he sells. Let P represent his total pay (in dollars), and let N represent the number of copies of History is Fun he sells. Write an equation relating P to N . Then use this equation to find his total pay if he sells 27 copies of History is Fun.

OpenStudy (anonymous):

P=2300+n120 So for n=27 P=2300+120x27 P=2300+3240 P=5540